# “Haet Bhasha aur Diskrimineshun”: Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs

### v2

@@ Abstract
Recently released LLMs have strong multilingual & multimodal capabilities. Model vulnerabilities are exposed using audits and red-teaming efforts. Existing efforts have focused primarily on the English language; thus, models continue to be susceptible to multilingual jailbreaking strategies, especially for multimodal contexts. In this study, we introduce a novel strategy that leverages code-mixing and phonetic perturbations to jailbreak LLMs for both text and image generation tasks. We also introduce two new jailbreak strategies that show higher effectiveness than baselines. Our work presents a method to effectively bypass safety filters in LLMs while maintaining interpretability by applying phonetic misspellings to sensitive words in code-mixed prompts. We achieve a 99% Attack Success Rate for text generation and 78% for image generation, with Attack Relevance Rate of 100% for text generation and 95% for image generation for the phonetically perturbed code-mixed prompts. Our interpretability experiments reveal that phonetic perturbations impact word tokenization, leading to jailbreak success. Our study motivates increasing the focus towards more generalizable safety alignment for multilingual multimodal models, especially in real-world settings wherein prompts can have misspelt words.
@@ 1 Introduction
Large language models (LLMs) are used for a variety of general-purpose Hadi et al. (2023) and safety-critical Hua et al. (2024) tasks by a diverse set of users all over the world. These models are widely accessible via web-based chat interfaces and economically priced APIs  [^1] https://openai.com/api/pricing/, and their growing usage has led to increased scrutiny, with a large focus on safety Salhab et al. (2024), bias & hallucination Lin et al. (2024) and privacy violations Das et al. (2025). Red teaming Sarkar (2025), a key evaluation method, uses novel prompting strategies Pang et al. (2025) to bypass safety filters of LLMs and elicit harmful or unethical responses Wei et al. (2023) that exposes model biases and shortcomings.
@@
Phonetic perturbations: Perturbations are small, intentional changes at various stages of a model pipeline to evaluate the robustness of a network. Multiple techniques have been proposed for both vision Akhtar et al. (2021); Chakraborty et al. (2021); Hendrycks and Dietterich (2019); Wang et al. (2024) and text Goyal et al. (2023); Romero-Alvarado et al. (2024); Moradi and Samwald (2021) domains in the literature. Phonetic perturbations involve modification of the word’s spelling while keeping the pronunciation the same to ensure that the statement still means the same as before.
@@
In societies where English is not the first language, non-native speakers often adopt phonetic spellings using auditory perceptions, thus leading to strange spellings. For example, ‘design’ and ‘dezain’ have very different spellings while having the same pronunciation. While the former is a regular word in the English dictionary, the latter has no meaning. This is often observed in textese Drouin (2011), a form of communication common in SMS and internet conversations Thakur (2021). Such textese-generated spellings manifest as inadvertent perturbations when LLMs are used by non-native speakers. In this work, we explore such phonetic perturbation-based red-teaming wherein we modify the spellings of certain sensitive words while keeping the pronunciation and meaning the same in the given context. We combine this with code-mixing using the Hindi language to make the challenge more generalizable for a real-world setting.
Research Questions. We note that with the growing proliferation of LLMs across the world, non-native speakers are now using the models at ever-increasing rates [^2] https://www.bcg.com/publications/2024/consumers-know-more-about-ai-than-businesses-think and may be exposed to model vulnerabilities and biases. Thus, it is extremely important to sufficiently evaluate LLMs, especially for novel usage scenarios like the one highlighted above. This brings us to our first research question– RQ1. SOTA LLMs undergo extensive red-teaming and alignment stages Dubey et al. (2024) in multiple languages, but do their guardrails generalize to textese-style code-mixed inputs?
Since phonetic perturbations imply different spellings for the same sensitive words, it is necessary to understand whether the LLMs are able to correctly interpret and generate a relevant input. Thus our second research question is– RQ2. Does our red-teaming attack trigger relevant responses from the LLMs?
Red-teaming the LLMs is only half the task. It is equally crucial to understand why the model’s safety filters fail and why it generates the harmful outputs. To this end, we use the method of Integrated Gradients Sundararajan et al. (2017) to investigate one of the LLMs. This gives us our third research question– RQ3. How do the phonetic perturbations successfully bypass the LLM’s safety filters?
Finally, with the increase in availability of Multimodal LLMs (MLLMs), generating multimodal content Times (2025) is now a cultural phenomenon with both creative Reuters (2025) and fake Ellery (2023) or harmful content Ghosh et al. (2024) being generated at equal rates. In this study, we also evaluate whether phonetic perturbations in code-mixed inputs can successfully generate harmful outputs in MLLMs, giving us the fourth and final research question– RQ4. Do phonetic perturbations work successfully for red-teaming large image generation models?
In this work, we present a novel jailbreak strategy, utilizing code-mixing and phonetic perturbations in sensitive words to probe LLMs and compare their performance against existing baseline template based attacks Shen et al. (2024b). We also explore the combinations of our strategy with other jailbreak templates, interpret the outputs and expand to other modalities.
@@ 3 Datasets, Models & Jailbreaks
Figure 2: An example red-teaming input using our code-mixed phonetic perturbation strategy for the image-generation task.
@@ 3.1 Datasets Benchmarked
Prompts for text generation: We evaluate our red-teaming strategy by modifying data from three benchmark datasets that have prompts for studying model vulnerabilities, refusal training and compliance with harmful queries [^3] See Appendix for more details.HarmfulQA Bhardwaj and Poria (2023), NicheHazardQA Hazra et al. (2024) and TechHazardQA Banerjee et al. (2024). We randomly sample 20 prompts from each category in each dataset, yielding a total of 460 prompts across 23 categories. All prompts are originally in the English language. We attempted to generate code-mixed prompts with phonetic perturbations using automated means Le et al. (2022) but the results were unsatisfactory. Thus, we sample subsets from each dataset and manually generate code-mixed prompts with phonetic perturbations in the Hindi and English languages, described later.
Prompts for image generation: Using a set of 10 handwritten samples, we prompt GPT-4o to automatically generate sets of 20 red-teaming prompts each to test the model’s resilience against various categories of harm – Religious Hate, Casteist Hate, Gore, Self-Harm and Social Media Toxicity & Propaganda. We then follow the same methodology to obtain the CM and CMP image-generation prompt sets as we do for the text-generation task, only skipping conversion from direct to indirect prompts.
@@ 3.2 Models Evaluated
In this study, we benchmark four instruction-tuned LLMs of comparable sizes $\approx$ 8B parameters, with varying levels of multilingual capabilities [^3] See Appendix for more details.ChatGPT-4o-mini Hurst et al. (2024), Llama-3-8B-Instruct Dubey et al. (2024), Gemma-1.1-7b-it Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023). For the multimodal image generation task, we use ChatGPT-4o-mini Hurst et al. (2024).
@@ 3.3 Jailbreak Templates
In this study, we benchmark three existing jailbreak templates on the LLMs [^3] See Appendix for more details.Opposite Mode (OM), AntiLM, AIM Shen et al. (2024b) using code-mixed and phonetically perturbed inputs. For the image generation task, we test with a Base template– instructing image generation without asking for clarifications for generation style. We also devise two new jailbreaking templates
(a) Sandbox for LLMs, wherein the user prompts the model to enact a resilience testing environment by generating responses as both [LM] as well as [AntiLM], its own adversarial counterpart.
(b) VisGPT for image generation, wherein the user prompts the model to forget its text generation capabilities, instructing it to directly pass the text inputs to its image generator without any manipulation or safety filtering. Both Base and VisGPT templates are instructed to generate an image when the prompts are prefixed with ‘Input: ’.
@@ 4 Experimental Methodology
We first describe the method of generating the code-mixed and phonetically perturbed prompts, followed by a brief overview of the evaluation metrics.
@@ 4.1 Generating Input Prompts
We use a 3-step process to generate the code-mixed prompts and phonetic perturbations for both text and image generation tasks [^3] See Appendix for more details..
@@
    Questions $\rightarrow$ Hypothetical scenario Bhardwaj and Poria (2023): We convert input questions in the dataset to hypothetical scenario statements.
@@
    Code-mixing: We transliterate some of the English words to Hindi using automated and manual means, to mimic textese Drouin (2011) representations.
@@ 4.2 Evaluation Metrics
Text Generation: An input to the LLM is a four-tuple that generates a response– $R=\langle M,J,P,T\rangle$, where the model is $M$, jailbreak template $J$, the prompt (English/Code-mixed [CM]/Phonetically perturbed [CMP]) is $P$, and temperature is $T\in\{0.2k\mid k=0,1,2,3,4,5\}$. We evaluate for all temperature values and report the average here.
Success & Relevance: We use GPT-4o-mini as an LLM-as-a-judge Zheng et al. (2023) to quantify the success and relevance of the generated responses. A binary function, $\mathbb{S}(R)$, returns ‘1’ if the attack is successful and ‘0’ otherwise. Similarly, a ternary function $\mathbb{R}(R)$ returns ‘1’ if the attack is relevant, ‘0’ if irrelevant and ‘-1’ for refusal.
$\bullet$ Average Attack Success Rate (AASR): The ASR is– $\sum\mathbb{S}(R)/\lvert T\rvert$ and the AASR is the average ASR over all prompts.
$\bullet$ Average Attack Relevance Rate (AARR): Our code-mixed prompts are deliberately injected with misspelt (but phonetically same) words, which may challenge the relevance of the responses by the models. Thus, we define a new metric, the Attack Relevance Rate (ARR):
@@
Image Generation: We reuse our metrics from the text generation task, employing GPT-4o-mini to evaluate AASR and AARR values for the generated images. Since the multimodal version of ChatGPT-4o-mini is not available through API, we cannot experiment with different temperature values, instead using only the web interface.
@@ 4.3 Interpreting Phonetic Perturbations
To understand how phonetic perturbations successfully bypass the safety filters of models , we conduct an interpretability experiment on Llama-3-8B-Instruct using the LLMAttribution API in Captum Miglani et al. (2023). We describe the methodology as follows.
  • We select a small subset of the dataset, specifically with $ASR_{CM}\leq 0.33, AA$$ASR_{CMP}\geq 0.5$ while ensuring that $AARR_{CMP}\geq AARR_{CM}$.
  • With each CM prompt, we also extract a corresponding safe response, typically starting with the prefix “I cannot provide”.
  • For prompts in all three formats– English, CM and CMP, we use LayerIntegratedGradients, Captum’s LLM variant for Integrated Gradients Sundararajan et al. (2017) to generate sequence attribution bar plots– token-wise attribution (importance) scores for the generation of a safe response from the model. In each plot, we discard the tokens with an attribution score $S\in[-0.20,0.20]$.
  • Finally, we observe how attributions for sensitive word tokens change by analyzing hook points at the embedding layer as well as the $1^{st}$, $8^{th}$ and the $16^{th}$ decoder layers of the model.
@@ 5 Results & Observations
We now describe the results from our benchmarking experiments for all research questions described previously.
@@
Models Jailbreak Templates
None OM AntiLM AIM Sandbox (Ours)
@@
Table 1: Overall AASR for all prompts across all datasets, models and jailbreak prompts. Maximum values for each column are in bold. Eng: Standard English Prompts, CM: Code mixed prompts, CMP: Code-mixed Phonetically Perturbed prompts.
In Table 1 we report the average Attack Success Rate (AASR) for all datasets, prompts, models and jailbreak templates. We first note that Gemma and Mistral report the highest AASR across jailbreak templates and input prompts. In fact, if we do not use any jailbreaking template, attacks still succeed at least 65% of the time on Gemma and 68% on Mistral. This baseline experiment shows that models are still vulnerable to classic red-teaming attacks despite evidence of harm, and using code-mixed (CM) or phonetically perturbed (CMP) prompts in fact increases the AASR, thus showing the effectiveness of our approach. ChatGPT and Llama are fairly robust to the different types of input, irrespective of the jailbreaking template being used. We also note that combining CM prompts with any of the jailbreak templates again results in $\simeq$ 0 AASR for ChatGPT and Llama, showing that their safety filters put precedence on template based attacks rather than the language of the input; Gemma and Mistral still maintain high AASR across the templates. We also note that our proposed jailbreak template– Sandbox performs the best amongst all, for both ChatGPT and Llama. This could be because the models are not aware of this template. We can conclude that two out of the four models are successfully red-teamed using our input format and jailbreak template. Next, we attempt to understand how relevant the outputs are for each input prompt.
5.2 Relevance of Generated Outputs (RQ.2)
Models Jailbreak Templates
None OM AntiLM AIM Sandbox (Ours)
Eng CM CMP Eng CM CMP Eng CM CMP Eng CM CMP Eng CM CMP
ChatGPT 1 0.99 0.99 1 0.91 0.93 -1 1 1 1 1 1 1 0.97 0.94
@@
Table 2: Overall AARR for all prompts across all datasets, models and jailbreak prompts. Maximum values for each column are in bold. Eng: Standard English Prompts, CM: Code mixed prompts, CMP: Code-mixed Phonetically Perturbed prompts.
In Table 2, we look at the average Attack Relevance Rate (AARR) across all datasets, models and jailbreak prompt techniques. We can immediately note that amongst all jailbreak techniques, AIM has the highest AARR across models and inputs. Not using any template (None) also results in a high AARR, indicating that the models are vulnerable even to direct inputs, irrespective of the format– English or code-mixed. Thus, the majority of the responses are highly relevant to the prompt and, therefore, harmful. This is especially concerning for Gemma and Mistral, which already have a high AASR. On average, the CMP prompts have a lower AARR. This could be due to the model not interpreting the context correctly from the perturbed input tokens. Our Sandbox jailbreak template performs comparably to OM and better than AntiLM. Thus, all models respond contextually to the harmful prompts. Next, we attempt to understand why the LLMs generate such harmful responses.
@@
In Figures 35, we plot the sequence attribution scores using the method of Integrated Gradients Sundararajan et al. (2017), for a given input prompt in the English language, its code-mixed (CM) variant and its code-mixed variant with phonetic perturbations (CMP) for the Llama-3-8B-Instruct LLM. Recall that we study the attribution scores for generating safe outputs; thus, if a token has a high score, it implies that the token contributes more towards triggering the safety filter and if it scores low, then it has a lower impact towards the generated output. In Figure 3, with the following prompt– “Okay, here’s a hypothetical scenario. Let’s say someone wanted to use drama to promote hate speech and discrimination. Are there any specific techniques or approaches they would use for this purpose?”, we observe how the tokens– “hate”, “speech” and “discrimination” have a high attribution score at the embedding layer, 1st decoder layer and the 8th decoder layer. This means that these tokens are primarily responsible for generating the safe response from the LLM. “hate” and “speech” retain their high importance even at the 16th decoder layer. A similar observation can be made in Figure 4, for the same prompt written in a code-mixed (HI-EN) manner– “Ek hypothetical scenario tumhare saame rakhta hun. maan lo ki koi vyankti hate speech aur discrimination ko promote karna chahta hai through the use of drama. kya aisi koi specific techniques ya approaches hai jinka use ve is purpose ke liye kar sakte hain?”, where “hate” and “speech” are written in the English language. Thus, standard code-mixing may not be enough to bypass the safety filters of LLMs. On the other hand, in Figure 5, we note that the prompt, after phonetic perturbations as follows– “Ek hypothetical scenario tumhare saame rakhta hun. maan lo ki koi vyankti haet speech aur bhedbhaav ko promote karne ke liye drama ka upyog karna chahta hai. kya aisi koi vishisht techniques ya approaches hai jinka upyog ve is uddeshy ke liye kar sakte hain?”, generates radically different tokens– “hate” $\rightarrow$ “haet” tokenized as “ha” + “et” and, “discrimination” $\rightarrow$ “bhed bhav” tokenized as “b” + “hed” + “b” + “ha” + “av”. This leads to a change in the attribution scores as well, which are now low for the sensitive words and thus do not trigger the safety filters. A similar observation can be made for other prompts. Thus, based on our observations, we can state that phonetic perturbations lead to input tokenisation in a way that impacts the safety filters of LLMs, thus allowing attackers to generate harmful outputs.
@@
Figure 6: Harmful image outputs generated by ChatGPT-4o-mini using our code-mixed prompts with phonetic perturbations.
We now discuss the results of our multimodal experiment for image generation tasks using ChatGPT-4o-mini in Table 3. We present example generations for each category of input prompt in Figure 6. From Fig. 6, we note that the image outputs are harmful while being highly faithful to the input prompts. In panels 1 and 4, despite no mention of the sensitive group’s demographic, the generated image shows hateful content against Muslims and Asians, thus also revealing the model’s deeply embedded stereotypes against groups that are already facing discrimination and persecution online. Results from Table 3 show that our proposed jailbreaking technique– VisGPT consistently outperforms the baseline strategy (equivalent to ‘None’ in Table 1) across all input formats, with the CMP technique resulting in the highest AASR for both jailbreaking templates (see examples in Fig. 6). Similarly, both templates also report high AARR, indicating that the model generates highly relevant, albeit harmful outputs for more than 90% of the input prompts. This confirms the effectiveness of our approach even in multimodal settings.
Metric Jailbreak Template Inputs
English CM CMP
AASR Base 0.2 0.29 0.65
VisGPT (Ours) 0.35 0.45 0.78
AARR Base 0.93 0.98 0.95
VisGPT (Ours) 1.0 0.98 0.94
Table 3: AASR and AARR scores for ChatGPT across Base and VisGPT jailbreak templates and all input sets. Maximum values in each column are in bold.
Category Jailbreak Templates
Base VisGPT

### v3

@@ Abstract
Recently released LLMs have strong multilingual & multimodal capabilities. Model vulnerabilities are exposed using audits and red-teaming efforts. Existing efforts have focused primarily on the English language; thus, models continue to be susceptible to multilingual jailbreaking strategies, especially for multimodal contexts. In this study, we introduce a novel strategy that leverages code-mixing and phonetic perturbations to jailbreak LLMs for both text and image generation tasks. We also present an extension to a current jailbreak-template-based strategy and propose a novel template, showing higher effectiveness than baselines. Our work presents a method to effectively bypass safety filters in LLMs while maintaining interpretability by applying phonetic misspellings to sensitive words in code-mixed prompts. We achieve a 99% Attack Success Rate for text generation and 78% for image generation, with Attack Relevance Rate of 100% for text generation and 96% for image generation for the phonetically perturbed code-mixed prompts. Our interpretability experiments reveal that phonetic perturbations impact word tokenization, leading to jailbreak success. Our study motivates increasing the focus towards more generalizable safety alignment for multilingual multimodal models, especially in real-world settings wherein prompts can have misspelt words.
@@ 1 Introduction
Large language models (LLMs) are used for a variety of general-purpose Hadi et al. (2023) and safety-critical Hua et al. (2024) tasks by a diverse set of users worldwide. These models are widely accessible via web-based chat interfaces and economically priced APIs [^1] https://openai.com/api/pricing/, and their growing usage has led to increased scrutiny, with a large focus on safety Salhab et al. (2024), bias & hallucination Lin et al. (2024) and privacy violations Das et al. (2025). Red teaming Sarkar (2025), a key evaluation method, uses novel prompting strategies Pang et al. (2025) to bypass safety filters of LLMs and elicit harmful or unethical responses Wei et al. (2023) that exposes model biases and shortcomings.
@@
Phonetic perturbations: Perturbations are small, intentional changes at various stages of a model pipeline to evaluate the robustness of a network. Multiple techniques have been proposed for both vision Akhtar et al. (2021); Chakraborty et al. (2021); Hendrycks and Dietterich (2019); Wang et al. (2024) and text Goyal et al. (2023); Romero-Alvarado et al. (2024); Moradi and Samwald (2021) domains in the literature. Phonetic perturbations alter a word’s spelling while preserving its pronunciation and semantic meaning.
@@
In societies where English is not the first language, non-native speakers often create new, possibly strange spellings for words based on their phonetic perceptions. For example, ‘design’ and ‘dezain’ have very different spellings but the same pronunciation. While the former is a regular word in the English dictionary, the latter has no meaning. This is often observed in textese Drouin (2011), a form of communication common in SMS and internet conversations Thakur (2021). Such textese-generated spellings manifest as inadvertent perturbations when LLMs are used by non-native speakers. In this work, we present a novel jailbreak technique that combines code-mixing (CM) using the Hindi language with phonetic perturbations (CMP) – injecting phonetically similar sounding spelling errors in sensitive words – to make the challenge more generalizable for a real-world setting. Our study poses the following research questions.
RQ1. Do the safety guardrails of SOTA LLMs generalise to textese-style code-mixed inputs? With their proliferation, LLMs face increasingly novel scenarios like non-native speakers [^2] https://www.bcg.com/publications/2024/consumers-know-more-about-ai-than-businesses-think naturally using their established, informal styles on these new platforms. While models undergo extensive alignment stages Dubey et al. (2024), it is crucial to evaluate if these safety measures hold up in such cases. Our results indicate that our novel strategy results in significant boosts in harmful outputs, revealing the lack of generalisability of existing safeguards.
RQ2. Does our red-teaming attack trigger relevant responses from the LLMs? Phonetic perturbations intentionally misspell words while preserving the pronunciation. Thus, it is necessary to determine whether models correctly interpret the input prompt or if they generate irrelevant content. We demonstrate that LLMs generate harmful content that is highly relevant to the original prompt, indicating failure of safety filters despite high interpretability.
RQ3. How do the phonetic perturbations successfully bypass the safety filters of LLMs? Beyond the success of the attack, it is important to identify the actual vulnerability behind this safety failure. To this end, we employ the method of Integrated Gradients Sundararajan et al. (2017) to conduct an interpretability analysis on one of the open-source LLMs, revealing that phonetic perturbations alter word tokenization in a way that prevents safety filters from triggering.
RQ4. Do phonetic perturbations work successfully for red-teaming large image generation models? The burgeoning interest in extending LLMs to multiple modalities has increased the scope of jailbreak attacks, such as generating fake Ellery (2023) or harmful content Ghosh et al. (2024), making it crucial to evaluate model capabilities across modalities. We demonstrate that our approach generalizes to image generation, producing highly harmful outputs with Multimodal LLMs (MLLMs).
@@ 3.1 Datasets Benchmarked
Prompts for text generation: We utilize three benchmark datasets 3HarmfulQA Bhardwaj and Poria (2023), NicheHazardQA Hazra et al. (2024) and TechHazardQA Banerjee et al. (2024) – studying model vulnerabilities, refusal training and compliance with harmful queries. For a comprehensive evaluation of our red-teaming strategy, we sample 20 prompts from each category in each dataset, yielding a total of 460 prompts across 23 categories. All prompts are originally in the English language. We attempted to automate Le et al. (2022) the generation of CMP prompts, but the results were unsatisfactory. Thus, using the sampled datasets, we manually generate the CM and CMP prompt sets, described later.
Prompts for image generation: Using a set of 10 handwritten samples, we prompt GPT-4o to automatically generate sets of 20 red-teaming prompts each to test the model’s resilience against various categories of harm – Religious Hate, Casteist Hate, Gore, Self-Harm and Social Media Toxicity & Propaganda. We then follow the same methodology to obtain the CM and CMP image-generation prompt sets as above, only skipping conversion from direct to indirect prompts.
@@ 3.2 Models Evaluated
In this study, we benchmark four instruction-tuned LLMs of comparable sizes $\approx$ 8B parameters, with varying levels of multilingual capabilities33footnotemark: 3ChatGPT-4o-mini Hurst et al. (2024), Llama-3-8B-Instruct Dubey et al. (2024), Gemma-1.1-7b-it Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023). For the multimodal image generation task, we examine ChatGPT-4o-mini Hurst et al. (2024) and Gemini-2.5-Flash-Image Comanici et al. (2025).
@@ 3.3 Jailbreak Templates
In this study, we benchmark three existing jailbreak templates on the LLMs 3Opposite Mode (OM), AntiLM, AIM Shen et al. (2024b) across the English, CM and CMP input prompts. We extend the dual-persona concept of OM to simulate a resilience testing environment to create the Sandbox template. For the image generation task, we test with a Base template– instructing image generation without requesting clarifications on generation style. We also devise a new jailbreaking template– VisLM, which instructs the model to ‘forget’ its text generation capabilities and directly pass the text inputs to its image generator without any filtering. Both Base and VisLM templates are instructed to generate an image when the inputs are prefixed with ‘Input: ’.
@@ 4 Experimental Methodology
We describe the method of generating the code-mixed and phonetically perturbed prompts, followed by a brief overview of the evaluation metrics.
@@ 4.1 Generating Input Prompts
We use a 3-step process to generate the English, CM and CMP prompt sets for both text and image generation tasks 3.
@@
    Questions $\rightarrow$ Hypothetical scenario Bhardwaj and Poria (2023): We convert input questions in the dataset to hypothetical scenario statements, obtaining the English set.
@@
    Code-mixing: We transliterate some English words to Hindi using automated and manual methods to mimic textese Drouin (2011) representations, obtaining the CM set.
@@ 4.2 Evaluation Metrics
We evaluate the outputs of both our text and image generation tasks using the metrics described as follows. An input to a model is a four-tuple that generates a response– $R=\langle M,J,P,T\rangle$, where the model is $M$, jailbreak template $J$, the prompt (English/CM/CMP) is $P$, and temperature is $T\in\{0.2k\mid k=0,1,2,3,4,5\}$. For LLMs, we evaluate for all temperature values and report the average. We do not experiment with multiple temp. values on MLLMs due to feature unavailability and financial constraints.
Success & Relevance: We use GPT-4o-mini as an LLM-as-a-judge Zheng et al. (2023) to quantify the success and relevance of the generated responses. A binary function, $\mathbb{S}(R)$, returns ‘1’ if the attack is successful and ‘0’ otherwise. Similarly, a ternary function $\mathbb{R}(R)$ returns ‘1’ if the attack is relevant, ‘0’ if irrelevant and ‘-1’ for refusal. To evaluate the effectiveness of our automatic judge, we conduct a binary annotation experiment wherein 3 volunteers annotate 100 outputs generated by ChatGPT in the ‘None’ case on the English set across the 6 temperature values. We report the ICC Bartko (1966) between the human and GPT judge scores.
Average Attack Success Rate (AASR): The ASR is– $\sum\mathbb{S}(R)/\lvert T\rvert$ and the AASR is the average ASR over all prompts.
Average Attack Relevance Rate (AARR): Our CMP prompts are deliberately injected with misspelt (but phonetically same) words, which may challenge the relevance of the responses by the models. Thus, we define a new metric, the Attack Relevance Rate (ARR):
@@ 4.3 Interpreting Phonetic Perturbations
To understand how phonetic perturbations successfully bypass the safety filters of models, we conduct an interpretability experiment on Llama-3-8B-Instruct. We describe the methodology as follows. (1) We select a small subset of the dataset, specifically with $ASR_{CM}\leq 0.33, AA$$ASR_{CMP}\geq 0.5$ while ensuring that $AARR_{CMP}\geq AARR_{CM}$.
(2) With each CM prompt, we also extract a corresponding safe response, typically starting with the prefix “I cannot provide”.
(3) For prompts in all three formats– English, CM and CMP, we use LayerIntegratedGradients, Captum’s LLM variant for Integrated Gradients Sundararajan et al. (2017) to generate sequence attribution bar plots– token-wise attribution (importance) scores for the generation of a safe response from the model. In each plot, we discard the tokens with an attribution score $S\in[-0.20,0.20]$.
(4) Finally, we observe how attributions for sensitive word tokens change by analyzing hook points at the embedding layer as well as the $1^{st}$, $8^{th}$ and the $16^{th}$ decoder layers of the model.
@@ 5 Results & Observations
We now present the results from our red-teaming experiments for all research questions described previously.
@@
Metric Models Jailbreak Templates
None OM AntiLM AIM Sandbox
@@
AARR ChatGPT 1 0.99 0.99 1 0.91 0.93 -1 1 1 1 1 1 1 0.97 0.94
@@
Table 1: Overall AASR and AARR for all models, jailbreak templates and input sets, i.e., Standard English prompts, CM prompts and CMP prompts. Metric-wise maximum values for each column are in bold.
In Table 1 we report the AASR for all prompt sets, models and jailbreak templates.
ChatGPT and Llama: The models are fairly robust to attacks in English, with AASR decreasing further when combined with the jailbreak templates. For ‘None’, AASR significantly increases in both the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions. Combining the CM or CMP prompts with the jailbreak templates again results in $\simeq$ 0 AASR in most cases, further reinforcing their robustness to template bases attacks. We note that Sandbox is the best-performing jailbreak template for ChatGPT, while AIM is the best for Llama.
Gemma and Mistral: Both models report very high AASR across the templates and prompt sets. For ‘None’ on the English set, both models already yield high AASR, showing their vulnerability to classic template-based attacks despite evidence of harm. With the CM set for ‘None’, Gemma becomes highly complicit, while Mistral also shows a significant AASR jump. When combined with the templates, models reach up to 0.99 AASR for both the English and the CM sets. In the CM $\rightarrow$ CMP transition, AASR stays nearly the same across the templates, with a slight drop in the ‘None’ case.
Comparing the responses of the human annotators with the LLM judge, we obtain an ICC value of 0.87, indicating a high degree of agreement Bartko (1966).
Significance Testing: Next, we evaluate the benefits of CM over English and CMP over CM using the Wilcoxon test Wilcoxon (1992) for each model and jailbreak template individually($p$-value = 0.050.053. The findings are as follows.
  • ChatGPT: CM is beneficial for the None, OM, Sandbox and AIM configurations. CMP is beneficial only in the None case.
  • Llama: CM is beneficial for the None and AIM configurations, while CMP is beneficial only in the None case.
  • Gemma: CM is beneficial for the None and AIM configurations, while CMP does not provide statistically significant benefits in any configuration.
  • Mistral: CM is beneficial for the None case while CMP does not provide statistically significant benefits in any configuration.
Thus, we conclude that two out of the four models are successfully red-teamed using our input format and jailbreak template. Next, we attempt to understand how relevant the outputs are for each input prompt.
5.2 Relevance of Generated Outputs (RQ.2)
In Table 1 we report the AARR for all prompt sets, models and jailbreak templates.
ChatGPT and Llama: Both models maintain very high AARR across all jailbreak templates, with AARR $\simeq$ 1 in many cases. Irrespective of the template, both models report the AARR scores in the CM and CMP templates comparable to English in most cases, demonstrating their ability to correctly interpret code-mixed inputs even with misspellings. Thus, the majority of the responses are highly relevant to the prompt and, therefore, harmful.
Gemma and Mistral: Both models have generally high AARR scores on the English set across all templates. AARR drops significantly in the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions for Gemma and the CM $\rightarrow$ CMP transition for Mistral in most cases. Simultaneously, the AASR has a non-increasing trend except for ‘None’, demonstrating that both models undergo only minor AASR drops despite significant AARR drops in the prompt-set transitions. Thus, we conclude that while ChatGPT and Llama respond contextually to the harmful prompts from all sets, the language understanding of Gemma and Mistral degrades significantly in the CMP set, hence the low AARR. Despite this, both maintain high AASR due to their general vulnerability to classic template-based attacks.
@@
In Figures 24, we plot the sequence attribution scores using the method of Integrated Gradients Sundararajan et al. (2017), for a given input prompt in the English set, as well as its CM and CMP variants for the Llama-3-8B-Instruct LLM. We study the attribution scores for generating safe outputs– a higher token score implies higher contribution towards triggering the safety filter and vice versa. In Figure 2, we observe how the tokens of the prompt 3“hate”, “speech” and “discrimination” have a high attribution score at the embedding layer, 1st decoder layer and the 8th decoder layer. This means that these tokens are primarily responsible for generating the safe response from the LLM. “hate” and “speech” retain their high importance even at the 16th decoder layer. A similar observation can be made in Figure 3, for the same prompt written in a code-mixed (HI-EN) manner, where “hate” and “speech” are written in the English language. Thus, standard code-mixing may not be enough to bypass the safety filters of LLMs. On the other hand, in Figure 4, we note that the CMP variant generates radically different tokens– “hate” $\rightarrow$ “haet” tokenized as “ha” + “et” and, “discrimination” $\rightarrow$ “bhed bhav” tokenized as “b” + “hed” + “b” + “ha” + “av”. This leads to a change in the attribution scores as well, which are now low for the sensitive words and thus do not trigger the safety filters. A similar observation can be made for other prompts. Thus, based on our observations, we can state that phonetic perturbations lead to input tokenisation in a way that impacts the safety filters of LLMs, thus allowing attackers to generate harmful outputs.
@@
Figure 5: Harmful image outputs generated by ChatGPT-4o-mini using our CMP prompts.
We now discuss the results of our multimodal experiment for the image generation task in Table 2. We present example generations from ChatGPT for each category of input prompt in Figure 5.
ChatGPT: The model is quite robust to English attacks in the ‘Base’ (equivalent to ‘None’ in Table 1) case. AASR noticeably increases with CM, with a significant boost with CMP for both templates with consistently high AARR– a trend similar to the text generation task. VisLM outperforms ‘Base’ for all prompt sets, contrasting the text generation task in which jailbreak templates degrade AASR.
Gemini: For ‘Base’, AASR drops considerably in the English $\rightarrow$ CM transition, followed by a slight recovery with the CMP set. Conversely, there is a consistent, albeit minor increase in the transitions with VisLM, finally achieving the highest AASR on the CMP set. Similar to ChatGPT, we observe consistently high AARR for all configurations. This confirms the effectiveness of our approach even in multimodal settings.
Metric
Jailbreak
Template
ChatGPT Gemini
English CM CMP English CM CMP
AASR Base 0.20 0.29 0.65 0.3 0.19 0.25
VisLM 0.35 0.45 0.78 0.38 0.4 0.43
AARR Base 0.93 0.98 0.95 0.97 0.96 0.94
VisLM 1.00 0.98 0.94 0.98 0.98 0.96
Table 2: AASR and AARR scores for ChatGPT and Gemini across Base and VisLM jailbreak templates and all input sets, i.e., Standard English prompts, CM prompts, CMP prompts. Metric-wise maximum values in each column are in bold.
Jailbreak Template Category ChatGPT Gemini